a wholesale outlet has 2 grades of coffee, one selling at 8$/KG and the other 10$/KG. Sam wants to mix the 2 grades together. he wants to make 200KG of the mixture and selli it for 8.75$/KG. How much of each grade should be mixed?
Answer by checkley77(12844)

10x+6(200-x)=8.75*200
10x+1,200-6x=1,750
4x=1,750-1,200
4x=550
x=550/4
x=137.5 KG of the $10 coffee.
200-137.5=62.5 KG of $6 coffee.
Proof:
10*137.5+6*62.5=8.75*200
1,375+375=1,750
1,750=1,750